# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ V.4 Modeling atomistic surface diffusion under high electric fields
As it was discussed in Section V.3, the partial charge induced by an external electric field on surface atoms affects the thermodynamic equilibrium due to additional forces of electrostatic nature. In turn, this affects the thermally activated dynamics of atoms inducing biases towards stronger fields in random migration of surface atoms, which is known as surface diffusion.
@@
A kinetic Monte Carlo (kMC) method parameterized to describe specifically the hopping events for surface atoms was successfully used to simulate the surface diffusion [65, 67]. The parameterization of such a model is based on calculations of migration barriers that are functions of the local atomic environment. Enabling the effects of an electric field in these simulations is not straightforward, since the continuous electron density distributions must fit within the discrete realm of the kMC method. Applying the methodology described in Section V.3, one can use a single formula to modify the set of existing barriers introducing the effect of an electric field.
[…]
@@
The same technique with the fixed effect of electric field on the value of migration barriers was used in simulations of surface tip sharpening [66]. Sharp surface features with high aspect ratio are frequently assumed to explain localized FE currents from seemingly flat metal surfaces [76]. However, such tips were shown not to survive on a metal surface without an external electric field [64] due to the surface energy minimization principle. When an electric field is applied, the dynamics of surface atoms may change because of electric polarization around them. The stronger an atom stands out from the surface the higher its polarization becomes, see Figure 49.
[…]
@@ VI.1 Thermal effects in electron emitting tips
The observed high values of field enhancement factors [76] imply the existence of field emitting tips, which either exist prior to high electric field application or grow on the surface as a result of field-induced surface dynamics. It is natural to consider the intensively emitting tips as a potential source of neutrals, which, in combination with the field-emitted electrons, initiate the plasma that forms the conducting channels and eventually leads to the discharge. The heat generated in the tip by the strong field emission currents affects both the electron emission and the atomic dynamics in sharp tips. Hence, it is necessary to calculate the temperature distribution in a tip of a given geometry. This problem is already quite complex and multi-physics in nature since it requires calculating the electric field distribution around the tip, the field emitted current on its surface, the internal distribution of the current and heat, and the solution of the heat equation for the temperature distribution, as summarized graphically in Fig. 57. The problem can be approached numerically in various levels of approximation. The numerical solution of the Laplace equation for the field distribution can be combined with the field emission and Ohm’s laws to estimate the generated heat in a protrusion of given geometry [69, 74, 118]. This approach provides an assessment of the Joule heating effect. However, any possible shape modification of the protrusion remains inaccessible.
@@
A dominating contribution of Joule heating can be assumed only for field emitters with apex curvature radius in the order of a m and higher, since tips of this size will heat up more efficiently through the processes which take place in the bulk. For nanometer-sharp surface features with a very large surface-to-volume ratio, the Nottingham effect [129, 130] becomes a leading process in heat exchange between the tip and the electrons leaving it [118, 14]. Furthermore, the increase of temperature within the electron-emitting tip results in the transition from cold field emission to thermionic emission, as well as the intermediate thermal-field regime, which needs to be described accurately in a single model.
@@
It is evident from the above that building a model that accurately captures the evolution of intensively emitting nanotips at elevated temperatures is extremely challenging and requires developing novel computational techniques. Such techniques evolved gradually during the past decade, starting with developing the first hybrid electrodynamics-molecular dynamics model ”HELMOD” [36]. The first attempt to introduce electron emission and thermal effects in HELMOD was performed by Parviainen et al. [135]. Eimre et al. [42] introduced a more accurate emission model covering all emission regimes based on the General Thermal Field (GTF) equations [68], including the Nottingham effect, but only for static tips. The first model that accurately included all emission and thermal effects on a dynamically evolving tip was proposed by Kyritsakis et al. [86], revealing a thermal runaway mechanism, thanks to the introduction of a more versatile FEM-based technique for electrodynamics [172] that allows for a multiscale extension of the simulation domain, as well as a generalized numerical tool for electron emission calculations that covers accurately for all emission regimes [85]. This model was later enhanced with Particle-In-Cell capabilities by Veske et al. [174], allowing for the inclusion of space charge effects.
VI.2 Thermal runaway and atom emission as a precursor of vacuum arcing
@@
The full dynamics of the tip evolution due to the pre-breakdown strong electron emission currents can be described by multi-physics multiscale models [172, 86, 174]. In this model, the atomic dynamics are calculated explicitly within the top part of the tip, while both the electric field, current density, and temperature distributions are calculated within a finite element mesh that extends to the entire height of the tip and the flat surface surrounding it, see Figure 52.
@@
The developed concurrent multiscale model allows us to peek into the dynamics of the atoms under an externally applied electric field, taking into account both the field-induced stresses and the emission-induced thermal effects. These dynamics play a paramount role in determining the behavior of an intensively field emitting nanotip, as they contribute the major positive feedback in the thermal runaway mechanism [86]. Figure 53 demonstrates this, as it compares the evolution of the maximum temperature in a static (blue dashed line) versus a dynamic (red solid line) simulated tip. Although the two cases exhibit the same temperature evolution in the initial stages, the dynamic re-arrangement of atoms after the emission-induced melting results in a thermal runaway that causes the temperature spikes of figure 53. Under tensile stress due to the interaction of the electric field with the charged surface, the tip apex sharpens after melting. The local field and emitted current density increase within the sharpened apex, increasing the generated heat and forcing the atoms to break the atomic bonds and evaporate while promoting further sharpening in a positive feedback loop. In the assumption of a static tip, where no shape modification occurs, the temperature does not rise above a given saturated value (blue dashed line in Figure 53). Underestimation of temperature dynamics within the current emitting tip can compromise the analysis of processes that take place in the pre-breakdown stage of vacuum arc onset.
@@
Figure 54 shows the dynamic evolution of the tip shape. The plot shows the tip’s maximum height for two different initial tip geometries. The insets show the tip shape and temperature distribution (color coding) at specific frames (designated on the plot). The initial aspect ratio of both tip geometries was about 30, but two different shank openings at the bottom of the tip were simulated, with ”thin” and ”wide” referring to a characteristic shank radius of 17 nm and 54 nm respectively, compared to 3 nm at the apex of the tip. One can see that (c) and (d) correspond to the spike in height, while the shape of the tip is highly changed showing narrow necking with a strong temperature increase (scale in the color bar). This necking and the corresponding extremely highly local temperatures cause the detachment of entire metal droplets, resulting in the height rapidly decreasing, and the tip becoming blunt again. However, as it is still molten and hot, the thermal runaway mechanism restarts, leading to the next spike in height, in the case of the thin tip. In contrast, this does not occur in the wide geometry due to the more efficient heat conduction into the bulk with the broader opening at the bottom. Still, even in the wide tip case, some repeated dynamic still continues.
@@
Counting the number of atoms detached from the top of a tip per time, one can estimate the evaporation rate of neutrals with respect to the emission rate of electrons given by the emission currents. These estimations were consistent with the values predicted in Ref. [166] (see next section for details), where it was shown that the plasma discharge takes place if 15 neutral atoms are injected into the system for 1000 electrons. Hence these studies allow to conclude that the thermal runaway process in a metal tip acts a precursor to a full plasma onset.
@@
Finally, we should note that the results shown in figure 53 are taken from [86], while the ones of 54 are from [172]. The latter was run with a more advanced version of FEMOCS, which included PIC-MCC simulations that accurately calculate the space charge. In contrast, the earlier ones of [86] had an approximate correction for the space charge suppression of the emission. However, there is no significant difference between the results of the two methods regarding the main conclusions they lead to, as described above.
@@ VII Plasma ignition and evolution
One of the processes in vacuum arc initiation that is paramount to understand is the initiation of plasma. The plasma ignition dynamics are extremely hard to study experimentally, as the phenomenon’s spatial and temporal scales are extremely small (nm-ns). However, advanced modeling has given significant insights into vacuum arc plasma onset physics, employing mainly particle-in-cell plasma simulations.
The main challenge posed by the nature of the phenomenon for these simulations is the complex interaction of the plasma with the metal surface, which is the main determining factor of the plasma evolution. The following sections give an overview of the simulation models developed for arc ignition and the main results they offered.
@@ VII.1 Arc ignition models based on flat surface morphology
The first attempts to simulate the vacuum arc plasma initiation process were made by Timko et al. [165, 167] with the one-dimensional (1D) Particle-In-Cell with Monte Carlo Collisions (PIC-MCC) model “ArcPIC”. This model treated the cathode surface as a source of neutrals and electrons. The injection flux of electrons is defined by the cathode electric field via the Fowler-Nordheim equation to model field emission, assuming a certain field enhancement factor , which is removed after the evaporation of a certain amount of neutrals. The injection flux of neutrals is assumed to be proportional to the flux of electrons with a given flux ratio . This model was later expanded to a 2D-axisymmetric geometry [166], assuming that most of of the cathode surface has a small distributed field enhancement and in the center of the system there is a microscopic field enhancing feature with field enhancement factor .
Figure 56 shows the evolution of particle densities as simulated by the above ArcPIC model. We see that in the beginning of the simulation a strong electron beam appears near the center of the simulation (), accompanied by the evaporation of a strong neutral cloud. At about 1 ns, a significant amount of ions appear near the tip, as a result of electron impact ionization. This ion cloud gradually condenses and expands, eventually filling up the entire simulation box, developing into a full dense plasma. This plasma is accompanied by a potential sheath forming a strong electric field near the cathode. This field is necessary to both maintain a field emitted electron flow into the plasma, as well as the bombardment of the cathode with sufficiently energetic ions that sputter neutrals and maintain a neutral flow.
@@ VII.2 Plasma ignition from intensively emitting protrusions
The quasi-planar models presented in section VII.1 (ArcPIC), in combination with the thermal runaway mechanisms revealed by the concurrent multi-physics approach presented in section VI offered significant qualitative insights into the plasma initiation mechanisms. Yet, they do not yet offer a precise quantitative description of the plasma initiation process that can make predictions comparable to experimental data. This is mainly because the low-scale FEMOCS simulations of section VI are disconnected from the larger-scale ArcPIC model. There is an indirect connection in that FEMOCS results justify the neutral injection assumptions of ArcPIC, but this does not capture the interplay of the plasma and tip runaway processes during tip thermal runaway. The concurrent coupling of PIC to the hybrid ED-MD model has been proposed and developed to capture this process.
The first step in this direction was performed by Veske et al. [174], which incorporated PIC-MCC routines in FEMOCS. A significant advancement of this compared to ArcPIC is that the FEMOCS PIC implementation is based on finite elements rather than ArcPIC’s finite differences. This is crucial as it allows unstructured meshes to capture vastly different spatial scales in the same simulations. Combined with a variable weight superparticle approach, this can yield a model that fully couples the surface and plasma dynamics.
@@
This model investigated the physical processes beyond the dynamic evolution of the tip that are most significant for the development of plasma, focusing on the plasma ignition itself, i.e., the mechanisms that produce and maintain the supply of vapor and the ones that ionize it. To this end, the full tip geometry is simulated, yet in a static manner. The evaporation rates on the surface of the tip are calculated based on the local temperatures and the corresponding PIC superparicles being injected accordingly. For the ionization, in addition to the standard impact ionization processes considered in ArcPIC, FEMOCS includes direct field ionization via tunneling, as well as double and triple-ionized species. Finally, unlike ArcPIC, FEMOCS considers a normal sputtering yield on the surface without introducing the empirical enhanced sputtering due to ion impact heating. This ion bombardment heating is formally taken into account in FEMOCS in the heating problem, as the energy of the impacting ions is introduced as a surface heat component in the heating problem. Finally, a similar circuit model to ArcPIC was considered, yet the implementation is much more general to potentially capture any linear coupling between the macroscopic electromagnetic power sources and the localized microscopic breakdown site (see sec. IX for details).
@@
Figure 58 shows the evolution of the temperature distribution on the tip along with the distribution of all superparticle species in the vacuum region. It is evident that a significant plasma starts on the side of the tip and then gradually expands around the surface area of the tip. The most significant insights of these simulations are shown in Figure 59, where we plot the evolution of the various processes occurring on the tip. At the beginning of the plasma process, we see that the dominant ionization mechanism is field ionization, which was not considered in previous models. The fact that FEMOCS resolves the local field distribution allows us to properly take into account this effect, as field ionization is significant only in the vicinity of the tip apex where the field strength is in the range of 10 GV/m. Yet, in later stages, when the plasma develops significantly and the field in the plasma core drops, the impact ionization becomes dominant, with the field ionization oscillating.
Finally, in the bottom graph of Figure 59, we see that the main source of neutrals in the system is thermal evaporation, with sputtering playing a minor role. The evaporation rate curve follows very closely and slightly exceeds the one by ion bombardment, which is expected considering the particle balance in the plasma system, with only a small excess of injected neutrals contributing to the growth of the plasma.

### v2

@@ V.4 Modeling atomistic surface diffusion under high electric fields
As discussed in Section V.3, the partial charge induced by an external electric field on surface atoms affects the thermodynamic equilibrium due to additional forces of an electrostatic nature. In turn, this affects the thermally activated dynamics of atoms, inducing biases towards stronger fields in the random migration of surface atoms, which is known as surface diffusion.
@@
A kinetic Monte Carlo (kMC) method that was parameterized to specifically describe the hopping events for surface atoms was successfully used to simulate surface diffusion [68, 70]. The parameterization of such a model is based on calculations of migration barriers that are functions of the local atomic environment. Enabling the effects of an electric field in these simulations is not straightforward since the continuous electron density distributions must fit within the discrete realm of the kMC method. Applying the methodology described in Section V.3, one can use a single formula to modify the set of existing barriers introducing the effect of an electric field.
Strictly speaking, the polarization characteristics around the surface atoms positioned in different atomic environments should differ, since the electron densities are also distributed differently around the atoms differently bound at the surface. Implementation of the environment-dependent electric field effects is a challenging task as the functional form becomes very complex and requires multiple calculations of these parameters for various atomic configurations at the surface. However, the use of fixed values for both characteristics in the KMC model of surface diffusion on the W tip under an electric field [69] already gave encouraging results which are shown in Fig. 51. In the figure, the left panels show the experimentally observed surface modification of the W tip during field emission (FE) experiments at elevated temperatures. Three rows in the experimental images (from the top down) show the FE mapping, the hard-sphere model tip reconstruction to match the corresponding FE maps above, and the corresponding scanning electron microscopy (SEM) images of the actual tip. Although there is a clear difference in length and time scales between the experiment and the simulation (m vs nm and s vs ns, respectively) the trend of surface modification is very similar. The kMC snapshots in the right panel of Fig. 51 show the evolution of a W tip. The top row shows the initial tip and the tip held at low electric fields, which do not induce sufficient bias in surface diffusion to influence the surface dynamics. The result is similar to the experiment on the left, where heating the tip leads to its facetting with a large fraction of the lowest energy surfaces. The increase in field promotes the growth of the surface at the top, since this allows maximal growth of the total length of sharp edges, where an electric field gradient appears giving rise to the biased diffusion. When the field becomes too large, the order gets confused, as any new defect on the surface will create the spot of the field gradient, reducing the order of the surface; see the last bottom image in the experimental and simulation panels.
@@
Generally, sharp surface features with high aspect ratio are often assumed to explain localized FE currents from seemingly flat metal surfaces [80]. However, such tips on a nanoscale were shown not to survive on a metal surface without an external electric field [67] due to the principle of minimization of surface energy. When an electric field is applied, the dynamics of surface atoms changes due to the redistribution of charge densities around them, as discussed in Section V.3. The stronger the atom stands out from the surface, the stronger the field enhancement and the higher the partial charge accumulates on it. The simulations where Eq. (6) was used to estimate the migration barriers[69] are shown in Fig. 52a. The top snapshot in Fig. 52a demonstrates an initial small surface feature, which develops into a sharp needle-shaped feature (bottom panel in Fig. 52b) due to the biased surface diffusion under an applied electric field.
The bias in surface diffusion caused by the presence of electric field gradients was also confirmed in the atomistic dynamic simulation augmented by collective variable hyperdynamics, which allowed a significant increase in the simulated time [83]. The snapshots in Fig. 52b show the calculations performed separately for different parts along the height of a surface tip. The three images stacked in Fig. 52b from bottom to top illustrate the atomic dynamics as a function of the surface electric field, which is shown by arrows colored according to the strength of the field. The color of the atoms indicates the magnitude of the partial charge: dark blue (close-to-zero charge) to bright red (0.5e). As one can see, at the bottom of the tip the field is low and no diffusion of surface atoms is observed. This changes significantly at the top of the tip where the field is the highest, and the atoms with the high partial charge move toward the stronger field more easily. Quantitatively, the biased effect in this work was assessed as an increase in drift velocity towards the top of the tip, in particular on the closed-packed surface. In these simulations, only the partial charge induced by the field near the surface atoms was estimated (see Section V.1) and included in simulations of surface dynamics instead of considering dipole moments, which was discussed in Section V.3. Following the drift velocity, the authors confirmed the effect of biased diffusion on metal surfaces under external high electric fields, particularly on close-packed surfaces where the energy barriers compare closer with the changes introduced by the interaction of surface charges with the external electric field.
@@ VI.1 Thermal effects in electron emitting tips
The observed high values of field enhancement factors [80] imply the existence of field emitting tips, which either exist prior to high electric field application or grow on the surface as a result of field-induced surface dynamics. It is natural to consider the intensively emitting tips as a potential source of neutrals, which, in combination with the field-emitted electrons, initiate the plasma that forms the conducting channels and eventually leads to the discharge. The heat generated in the tip by the strong field emission currents affects both the electron emission and the atomic dynamics in sharp tips. Hence, it is necessary to calculate the temperature distribution in a tip of a given geometry. This problem is already quite complex and multi-physics in nature since it requires calculating the electric field distribution around the tip, the field emitted current on its surface, the internal distribution of the current and heat, and the solution of the heat equation for the temperature distribution. The problem can be approached numerically in various levels of approximation. The numerical solution of the Laplace equation for the field distribution can be combined with the field emission and Ohm’s laws to estimate the generated heat in a protrusion of given geometry [72, 78, 127]. This approach provides an assessment of the Joule heating effect. However, any possible shape modification of the protrusion remains inaccessible.
@@
A dominating contribution of Joule heating can be assumed only for field emitters with an apex curvature radius on the order of a m and higher, since tips of this size will heat up more efficiently through the processes that take place in the bulk. For nanometer-sharp surface features with a very large surface-to-volume ratio, the Nottingham effect [138, 139] becomes a leading process in heat exchange between the tip and the electrons leaving it [127, 14]. Furthermore, the increase in temperature within the electron-emitting tip results in the transition from cold field emission to thermionic emission, as well as the intermediate thermal-field regime, which needs to be described accurately in a single model.
@@
It is evident from the above that building a model that accurately captures the evolution of intensively emitting nanotips at elevated temperatures is extremely challenging and requires developing novel computational techniques. Such techniques evolved gradually during the past decade, starting with the development of the first hybrid electrodynamics-molecular dynamics model HELMOD [39]. The first attempt to introduce electron emission and thermal effects in HELMOD was performed by [144]. [45] introduced a more accurate emission model that covers all emission regimes based on the General Thermal Field (GTF) equations [71], including the Nottingham effect, but only for static tips. The first model that accurately included all emission and thermal effects on a dynamically evolving tip was proposed by [91]. It revealed a thermal runaway mechanism, thanks to the introduction of a more versatile FEM-based technique for electrodynamics [182] that allows for a multiscale extension of the simulation domain, as well as a generalized numerical tool for electron emission calculations that accurately covers all emission regimes [90]. This model was later enhanced with Particle-In-Cell capabilities by [184], allowing the inclusion of space-charge effects.
VI.2 Thermal runaway and atom emission as a precursor of vacuum breakdown
@@
The full dynamics of the tip evolution due to the pre-breakdown strong electron emission currents can be described by multi-physics multiscale models [182, 91, 184]. In this model, the atomic dynamics is calculated explicitly within the top part of the tip, while both the electric field, current density, and temperature distributions are calculated within a finite element mesh that extends to the entire height of the tip and the flat surface surrounding it, see Fig. 53.
@@
The developed concurrent multiscale model allows us to peek into the dynamics of the atoms under an externally applied electric field, taking into account both the field-induced stresses and the emission-induced thermal effects. These dynamics play a paramount role in determining the behavior of an intense field-emitting nanotip, as they contribute the major positive feedback in the thermal runaway mechanism [91]. Figure 54 demonstrates this, comparing the evolution of the maximum temperature in a static (blue dashed line) to a dynamic (red solid line) simulated tip. Although the two cases exhibit the same temperature evolution in the initial stages, the dynamic rearrangement of atoms after the emission-induced melting results in a thermal runaway that causes the temperature spikes of Fig. 54. Under tensile stress due to the interaction of the electric field with the charged surface, the tip apex sharpens after melting. The local field and emitted current density increase within the sharpened apex, increasing the generated heat and forcing the atoms to break the atomic bonds and evaporate while promoting further sharpening in a positive feedback loop. In the assumption of a static tip, where no shape modification occurs, the temperature does not rise above a given saturated value (blue dashed line in Fig. 54). Underestimation of temperature dynamics within the current-emitting tip can compromise the analysis of processes that take place in the pre-breakdown stage of vacuum arc onset.
@@
Figure 55 shows the dynamic evolution of the shape of the tip. The plot shows the maximum height of the tip for two different initial tip geometries. The insets show the tip shape and temperature distribution (color coding) in specific frames (designated on the plot). The initial aspect ratio of both tip geometries was about 30, but two different shank openings at the bottom of the tip were simulated, with ”thin” and ”wide” referring to a characteristic shank radius of 17 nm and 54 nm respectively, compared to 3 nm at the apex of the tip. One can see that (c) and (d) correspond to the spike in height, while the shape of the tip is highly changed showing narrow necking with a strong temperature increase (scale in the color bar). This necking and the corresponding extremely highly local temperatures cause the detachment of entire metal droplets, resulting in the height rapidly decreasing, and the tip becoming blunt again. However, as it is still molten and hot, the thermal runaway mechanism restarts, leading to the next spike in height, in the case of the thin tip. In contrast, this does not occur in the wide geometry due to the more efficient heat conduction into the bulk with the broader opening at the bottom. Still, even in the wide tip case, some repeated dynamic still continues.
@@
By counting the number of atoms detached from the top of a tip per unit time, one can estimate the evaporation rate of neutrals with respect to the emission rate of electrons given by the emission currents. These estimations were consistent with the values predicted in Ref. [176] (see next section for details), where it was shown that the plasma discharge takes place if 15 neutral atoms are injected into the system for 1000 electrons. Hence these studies suggest that the thermal runaway process in a metal tip acts a precursor to a full plasma onset.
@@
Finally, we should note that the results shown in Fig. 54 are taken from [91], while the ones of Fig. 55 are from [182]. The latter was run with a more advanced version of FEMOCS, which included PIC-MCC simulations that accurately calculate the space charge. In contrast, the earlier ones of [91] had an approximate correction for the space charge suppression of the emission. However, there is no significant difference between the results of the two methods regarding the main conclusions they lead to, as described above.
@@ VII Plasma ignition and evolution
One of the processes in vacuum arc initiation that is paramount to understanding is the initiation of plasma. The plasma ignition dynamics are extremely hard to study experimentally, as the phenomenon’s spatial and temporal scales are extremely small (nm-ns). However, advanced modeling has given significant insights into vacuum-arc plasma-onset physics, employing mainly particle-in-cell plasma simulations.
The main challenge posed for these simulations is the complex interaction of the plasma with the metal surface, which is the main determining factor of plasma evolution. The following sections give an overview of the numerical simulation models developed for arc ignition and the main results they offered.
@@ VII.1 Arc ignition models based on flat surface morphology
Description of the arc ignition processes, from electron emission to plasma initiation, has been discussed semi-quantitatively by several authors in the past, see for example [26, 190, 119]. The advent of numerical simulations allowed for the research of a quantitative relationship between the field-emitted current and the amount of neutral atoms emitted required to establish a runaway plasma. However, these early attempts were based on a static effusion model of the injected neutrals [84] or simply a static background of neutrals [113]. The first attempt to simulate the vacuum arc plasma initiation process with a self-consistent relationship between electron emission and neutral injection was made by Timko et al. [175, 177] with the one-dimensional (1D) Particle-In-Cell with Monte Carlo Collisions (PIC-MCC) model “ArcPIC”. This model treated the cathode surface as a source of neutrals and electrons. The injection flux of electrons is defined by the cathode electric field via the Fowler-Nordheim equation to model field emission, assuming a certain field enhancement factor , which is removed after the evaporation of a certain amount of neutrals. The injection flux of neutrals is assumed to be proportional to the flux of electrons with a given flux ratio . This model was later expanded to a 2D axisymmetric geometry [176], assuming that most of the cathode surface has a small distributed field enhancement and in the center of the system there is a microscopic field enhancement feature with a field enhancement factor . A schematic representation of this model is shown in Fig. 56.
Figure 57 shows the evolution of particle densities as simulated by the above ArcPIC model. We see that in the beginning of the simulation a strong electron beam appears near the center of the simulation (), accompanied by the evaporation of a strong neutral cloud. At about 1 ns, a significant number of ions appear near the tip, as a result of electron impact ionization. This ion cloud gradually condenses and expands, eventually filling up the entire simulation box, developing into a full dense plasma. This plasma is accompanied by a potential sheath forming a strong electric field near the cathode. This field is necessary to maintain both a field-emitted electron flow into the plasma and the bombardment of the cathode with sufficiently energetic ions that sputter neutrals and maintain a neutral flow.
@@ VII.2 Plasma ignition from intensively emitting protrusions
The quasi-planar models presented in section VII.1 (ArcPIC), in combination with the thermal runaway mechanisms revealed by the concurrent multi-physics approach presented in section VI offered significant qualitative insights into the plasma initiation mechanisms. Yet, they do not yet offer a precise quantitative description of the plasma initiation process that can make predictions comparable to experimental data. This is mainly because the small-scale FEMOCS simulations of section VI are disconnected from the large-scale ArcPIC model. There is an indirect connection in that FEMOCS results justify the neutral injection assumptions of ArcPIC, but this does not capture the interplay of the plasma and tip runaway processes during tip thermal runaway. The concurrent coupling of PIC to the hybrid ED-MD model has been proposed and developed to capture this process.
The first step in this direction was performed by [184], which incorporated PIC-MCC routines in FEMOCS. A significant advancement of this compared to ArcPIC is that the FEMOCS PIC implementation is based on finite elements rather than ArcPIC’s finite differences. This is crucial as it allows unstructured meshes to capture vastly different spatial scales in the same simulations. Combined with a variable weight superparticle approach, this can yield a model that fully couples the surface and plasma dynamics.
@@
This model investigated the physical processes beyond the dynamic evolution of the tip that are most significant for the development of plasma, focusing on the plasma ignition itself, i.e., the mechanisms that produce and maintain the supply of vapor and the ones that ionize it. To this end, the full tip geometry is simulated, yet in a static manner. The evaporation rates on the surface of the tip are calculated based on the local temperatures and the corresponding PIC superparticles being injected accordingly. For the ionization, in addition to the standard impact ionization processes considered in ArcPIC, FEMOCS includes direct field ionization via tunneling, as well as double and triple-ionized species. Finally, unlike ArcPIC, FEMOCS considers a normal sputtering yield on the surface without introducing the empirical enhanced sputtering due to ion impact heating. This ion bombardment heating is formally taken into account in FEMOCS in the heating problem, as the energy of the impacting ions is introduced as a surface heat component in the heating problem. Finally, a similar circuit model to ArcPIC was considered, yet the implementation is much more general to potentially capture any linear coupling between the macroscopic electromagnetic power sources and the localized microscopic breakdown site (see Sec. IX for details).
@@
Figure 59 shows the evolution of the temperature distribution on the tip along with the distribution of all superparticle species in the vacuum region. It is evident that a significant plasma starts on the side of the tip and then gradually expands around the surface area of the tip. The most significant insights of these simulations are shown in Fig. 60, where we plot the evolution of the various processes occurring on the tip. At the beginning of the plasma process, we see that the dominant ionization mechanism is field ionization, which was not considered in previous models. The fact that FEMOCS resolves the local field distribution allows us to properly take into account this effect, as field ionization is significant only in the vicinity of the tip apex where the field strength is in the range of 10 GV/m. Yet, in later stages, when the plasma develops significantly and the field in the plasma core drops, the impact ionization becomes dominant, with the field ionization oscillating.
Finally, in the bottom graph of Fig. 60, we see that the main source of neutrals in the system is thermal evaporation, with sputtering playing a minor role. The evaporation rate curve follows very closely and slightly exceeds the one by ion bombardment, which is expected considering the particle balance in the plasma system, with only a small excess of injected neutrals contributing to the growth of the plasma.